Please help me solve this problem.
The measure of an angle is 2/3 the measure of its supplement. Find the measure
of the angle and its supplement.

Answer by jayanthihemkos@yahoo.com(87)

The measure of an angle is 2/3 the measure of its supplement. Find the measure
of the angle and its supplement.
let one angle be = x
then its supplement =( 180-x)
according to the sum we have
x = 2/3 (180-x)
x = 2/3*180 - 2/3*x
x = 2*60-2/3x
x+2/3*x= 120-2/3*x+2/3*x
3x+2x/3 = 120
5x/3*3= 120*3
5x = 360
5x/5 = 360/5
x = 72 degrees
the supplement angle = 180-72 = 108 degrees